HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

CHARACTERISATION OF THE GROUND AND LOW-LYING EXCITED STATES OF MgO^{+} BY PFI-ZEKE PHOTOELECTRON SPECTROSCOPY

\underline{JOEL R. SCHMITZ} , CARLA KREIS, FRÉDÉRIC MERKT, Department of Chemistry and Applied Biosciences, ETH Zürich, Zürich, Switzerland

We report on the characterisation of the rovibrational structure of the ground and first excited electronic states of MgO^{+} by high-resolution pulsed-field ionization zero-kinetic-energy (PFI-ZEKE) photoelectron spectroscopy. Rotationally cold (T_{rot} = 5K) MgO molecules in the X ^{1}Σ^{+} (v = 0 − 2) levels were generated in a supersonic expansion of a 0.1% N_{2}O:He gas mixture following laser ablation off a magnesium (Mg) rod [1]. The rovibrational ionization thresholds corresponding to both spin-orbit components (Ω = \frac{1}{2}, \frac{3}{2}) of the X^{+} ^{2}Π_{Ω} (v^{+} = 0 − 10) states as well as to the first excited A^{+} ^{2}Σ_{\frac{1}{2}} ^{+} (v^{+} = 0 − 10) state were reached in a resonant 1+1′ two-photon excitation sequence via the rovibrational levels of the F^{1}Π, E^{1}Σ^{+}, G^{1}Π and 3^{3}Π_{2} intermediate levels of MgO studied previously by Breckenridge and coworkers [2,3]. Our new results include accurate values for the adiabatic ionization energy of MgO and for the dissociation energies of the MgO X ^{1}Σ^{+} and MgO^{+} X^{+} 2Π_{Ω} and A^{+} ^{2}Σ+_{\frac{1}{2}} states. This work is carried out in the context of our studies of the rovibrational structure of doubly charged dications by high-resolution PFI-ZEKE spectroscopy of singly-charged cations following the approach recently taken to characterise the ground state of the thermodynamically stable dication MgAr^{2+} [4]. The talk will present a roadmap towards characterising the ground state of MgO^{2+} by resonant multiphoton excitation via electronically excited states of MgO^{+}. The experiments will reveal whether MgO^{2+} is thermodynamically stable as predicted in Ref. [5] or metastable as predicted in Ref. [6].

[1] T.D. Persinger, D.J. Frohman, W.M. Fawzy and M.C. Heaven, J. Phys. Chem. 153, 054308 (2020).

[2] J. Wang and W.H. Breckenridge, J. Chem. Phys. 124, 124308 (2006).

[3] D. Bellert, K.L. Burns, N.T. Van-Oanh, J. Wang and W. Breckenridge, Chem. Phys. Lett. 381, 381–384 (2003).

[4] D. Wehrli, M. Génévriez and F. Merkt, Phys. Chem. Chem. Phys. 23, 10978–10987 (2021).

[5] R. Linguerri, M. Hochlaf, M.C. Bacchus-Montabonel and M. Desouter-Lecomte, Phys. Chem. Chem. Phys. 15, 824–831 (2013).

[6] M. Kolbuszewski and J.S. Wright, Chem. Phys. Lett. 218, 338–342 (1994).

p-number: p079  Submitted on Mon, 12 Jun 2023 13:40:34 +0200